Given that p(x) =(k/2^x)s a probability distribution for a random variable that can take on the values x = 0, 1, 2, 3, and 4.

a. Find k.

b. Find the expression for the cdf F(x) of the random variable.


1
Expert's answer
2021-07-05T09:11:10-0400

a.


"p=\dfrac{k}{2^x}, x=0,1,2,3,4"

"\dfrac{k}{2^0}+\dfrac{k}{2^1}+\dfrac{k}{2^2}+\dfrac{k}{2^3}+\dfrac{k}{2^4}=1"

"\dfrac{k}{16}(16+8+4+2+1)=1"

"k=\dfrac{16}{31}"

b.


"F(x)=P(X\leq x)=\sum_{y:y\leq x}p(y)"

"F(x)=0, x<0"

"F(x)=16/31, 0\leq x<1"

"F(x)=16/31+8/31=24/31, 1\leq x<2"

"F(x)=24/31+4/31=28/31, 2\leq x<3"

"F(x)=28/31+2/31=30/31, 3\leq x<4"

"F(x)=30/31+1/31=1, x\geq4"

"F(x) = \begin{cases}
 0, &\text{for } x<0 \\
16/31, &\text{for } 0\leq x<1\\
24/31, &\text{for } 1\leq x<2\\
 28/31,&\text{for } 2\leq x<3\\
 30/31,&\text{for } 3\leq x<4\\
1,&\text{for } x\geq4.\\
\end{cases}"